the letters m,i,s,s,i,s,s,s,p,p,i,r,i,v,e,r are written on cards and placed face down. what is the probability that a random selection of one card will be an s?
1
Expert's answer
2012-07-20T07:25:25-0400
Here we have that all possibilities are equally likely. Hence
P(get s)= n(s)/16 = 5/16 .
& Here 16 stands for the total number of cards.